40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13984     Entry time: Mon Jun 18 19:47:02 2018
Author: gautam 
Type: Update 
Category: General 
Subject: MICH actuator calibration 

Summary:

The actuator (pendulum) gains for the Beam Splitter and the two ITMs were measured to be:

BS: 9.54 +/- 0.05 nm/ct [100 ohm series resistor in coil driver board]

ITMX: 2.44 +/- 0.01 nm/ct [400 ohm series resistor in coil driver board]

ITMY: 2.44 +/- 0.02 nm/ct [400 ohm series resistor in coil driver board]

Counts here refers to DAC counts at the output of the coil filter banks (as opposed to counts at the LSC servo output). The dominant (systematic) uncertainty (which isn't quoted here) in this measurement is the determination of the peak-to-peak swing of the dark port sensor, AS55_Q. I estimate this error to be ~1ct/33cts = 3%. These values are surprisingly consistent with one another once we take into account the series resistance.

Details:

The last time this was done, we used ASDC to do the measurement. But I don't know what signal conditioning ASDC undergoes (in PD or in readout electronics). In any case, in my early trials yesterday, ASDC was behaving unpredictably. So I decided to do redo the measurement.

[Attachment #1]- Flowchart describing the calibration procedure.

[Attachment #2] - AS55_Q output while the Michelson was freeswinging. I had first aligned the ITMs using ASS. The peak-to-peak value of this corresponds to \lambda/4. So we know AS55_Q in terms of cts/m of MICH displacement.

[Attachment #3] - Magnitudes of transfer function from moving one of the MICH optics, to the now calibrated AS55_Q. Fits are to a shape a/f^2, with a being the fitted parameter. Coherence during the measurement is also plotted.

  • Note that the excitation is applied to the channels C1:SUS-<optic>_LSC_EXC, for <optic> in [BS, ITMX, ITMY]. But since my de-whitening board re-work to remove the analog x3 gain, there is a digital x3 gain in the coil driver filter banks. So while the calibration numbers given above are accurate, be aware that when using them for sensing matrix measurements etc, you have to multiply these by x3.
  • Furthermore, moving the BS by x results in a Michelson length change of \sqrt{2}x, and this has been factored into the above number.

Next Steps:

  1. Now that I have a calibration I trust more, re-analyze my DRMI sensing matrix data. Actually the sensing response numbers aren't significantly different from what I have been assuming. It's just that in terms of counts applied at the LSC input of a suspension, there is a digital x3 gain that has to be explicitly factored in.
  2. Calibrate POX and POY by locking the arms and driving the now calibrated ITMs by a known number of counts.
  3. Calibrate the ETMs, and MC1/MC2/MC3 by looking at calibrated POX/POY.
  4. Lock DRMI, and calibrate SRM and PRM.

Reference:

[1] - http://www.phys.ufl.edu/~bernard/papers/CQG20_S903.pdf

Attachment 1: AS55cal_process.pdf  48 kB  Uploaded Mon Jun 18 23:55:44 2018  | Hide | Hide all | Show all
AS55cal_process.pdf
Attachment 2: AS55cal.pdf  81 kB  Uploaded Tue Jun 19 00:08:16 2018  | Show | Hide all | Show all
Attachment 3: MICH_act_calib.pdf  33 kB  Uploaded Tue Jun 19 00:08:22 2018  | Show | Hide all | Show all
ELOG V3.1.3-